OpenStudy (anonymous):

How does this model demonstrate the Pythagorean Theorem? A. The sum of the lengths of the shortest and the longest sides is equal to twice the length of the middle side. So double the length of the longer leg of any right triangle is equal to the sum of the shorter leg and the hypotenuse. B. The sum of the area of the two smaller squares is equal to the area of the larger square. So the sum of the lengths of the two legs of any right triangle squared is equal to the length of the hypotenuse squared.

OpenStudy (anonymous):

C. The sum of the area of the smallest and the largest squares is equal to the area of the middle square. So the sum of the lengths of the shorter leg and the hypotenuse of any right triangle squared is equal to the length of the middle leg squared. D. The length of the longest side minus two equals the length of the middle side. The length of the middle side minus two equals the length of the shortest side. So the length of the short leg of any right triangle is equal to the length of the middle leg minus 2, and the length of the hypotenuse is equal to the length of the middle leg plus 2.

OpenStudy (sweetburger):

Go with B. as if you add up the area of each shorter square its 64+36 = 100 which is the area of the larger square This is really hard to explain through OS but this fits directly into the Pythagorean theorem.

OpenStudy (sweetburger):

think about it like this 1 side 6^2 = 36 lets say that was a^2 and then 1 other side is 8^2 = 64 lets tsay that is b^2 and then the last side is 10^2= 100 and lets say that is c^2 so if you do this a^2 + b^2 = c^2 then 6^2 + 8^2 = 10^2

OpenStudy (sweetburger):

sorry that may be a bit confusing